ems TLN Group-Revenue by Division Year 1 (millilon) Year 2 (Total $5 milion) Advertising Branding 0.6 22% 32% 0.5 1.8 Public Relations 20% 1.1 26% In Year 3, Public Relations revenue (as a percentage of total revenue) increased by the same amount as the previous year, while total revenue increased by 20%, How much revenue did Public Relations make in Year 3? O $1.65 milion () $170 milion | | O $1.75 milion | ○ $1.80 milion。° Cannot Say Copyright IBM Corp. 2012, 2019 8
Answer #1

Total Revenue in year 2= =$5 million

Total Revenue in year3=5*(1+20%)=$6 million

Public Relation Revenue in Year 2=20% of Total Revenue

Public Relation Revenue in Year 1=$0.5 million

Public Relation Revenue as a percentage of total revenue in year 1=0.5/(0.6+0.5+1.1+1.8)=12.5%

Increase in Public Relation Revenue in Year 2 (compared to year 1)=20%-12.5%=7.5%

Public Relation revenue in year 3=20%+7.5%=27.5% of total revenue in year 3=6*27.5%=$1.65 million